OpenStudy (dallaslifebaby01):

What is the next term in the pattern 2, 9, 11, 20, 31, ...?

OpenStudy (freckles):

what is 2+9=? 11+20=?

OpenStudy (dallaslifebaby01):

11 and 32

OpenStudy (freckles):

11 and 31 right?

OpenStudy (dallaslifebaby01):

*31

OpenStudy (dan815):

i was gonna say i make these addition mistakes too some times :P

OpenStudy (freckles):

so can you tell what is going on in your sequence now?

OpenStudy (dallaslifebaby01):

yeah i get it know thanks

OpenStudy (freckles):

a very similar sequence is the fibnnoci sequence except it starts with 0 and1 0,1,1,2,3,5,8 like you are adding the previous two terms to find the next term after those two terms like 0+1=1 which is the third term 1+2=3 which is the fourth term and so on...

OpenStudy (freckles):

so what is the next term in your seqence

OpenStudy (dallaslifebaby01):

51

OpenStudy (freckles):

yes 20+31 which is 51 great job

OpenStudy (dallaslifebaby01):

thanks

OpenStudy (freckles):

and I skipped 9+11 earlier which is 20 but you seen the pattern I was getting at with the other two terms